Warephase Corporation has preferred stock outstanding. The stock has a 16% dividend rate. The stock’s market price is $80 per share, and its par value is $75. If new shares are issued, the firm will pay $3.50 per share in flotation costs. The corporate tax rate is 21%. What is the company’s cost of preferred stock financing?
Answer #1

Annual dividend=75*16%=$12

Hence cost of preferred stock=Annual dividend/(Current price-Flotation cost)

=12/(80-3.5)

=12/76.5

=15.69%(Approx)

NOTE:Cost of preferred stock is not affected by corporate tax rate.